OpenStudy (anonymous):

Convert 5x+y=2 from standard form to slope intercept form

OpenStudy (dtan5457):

slope intercept form= y=mx+b

OpenStudy (dtan5457):

subtract the 5x

OpenStudy (anonymous):

5x+y+5x=2+5x

OpenStudy (danjs):

you added 5x, try subtracting instead

OpenStudy (danjs):

5x - 5x = 0

OpenStudy (danjs):

5x - 5x + y = 2 - 5x . .

OpenStudy (anonymous):

5x-y-5x=2-5x

OpenStudy (danjs):

why did you change the +y to a -y?

OpenStudy (anonymous):

Oops sorry didn't mean to

OpenStudy (danjs):

5x + y = 2 (take away 5x from both sides) 5x - 5x + y = 2 - 5x y = -5x + 2

OpenStudy (danjs):

That is the slope-intercept form y = mx +b in this case the slope m = -5, and the y intercept b = 2

OpenStudy (anonymous):

Thank you

OpenStudy (danjs):

welcome